# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ Abstract
We introduce new invariants of the projective plane (and, more generally, of certain toric surfaces) that arise from appropriate enumeration of real curves of genus one and two. These invariants admit a refinement (according to the quantum index) similar to the one introduced by Grigory Mikhalkin in the rational case.
MSC-2010 classification: Primary 14N10, Secondary 14J26, 14P05, 14T90
@@ 1 Introduction
Refined enumerative geometry, initiated in [1, 7], became one of the central topics in enumerative geometry with important links to closed and open Gromov-Witten invariants and to Donaldson-Thomas invariants. In a big part of known examples, refined invariants appear as one-parameter deformations of complex enumerative invariants (see, for example, [1, 4, 6, 7]). In his groundbreaking paper [19], G. Mikhalkin has proposed a refined invariant provided by enumeration of real rational curves and has related this invariant to the refined tropical invariants of F. Block and L. Göttsche [1]. Namely, he introduced an integer-valued quantum index for real algebraic curves in toric surfaces. To have a quantum index, a real curve should satisfy certain assumptions: it has to intersect toric divisors only at real points and to be irreducible and separating; the latter condition means that, in the complex point set of the normalization of the curve, the complement of the real part is disconnected, i.e., formed by two halves exchanged by the complex conjugation (in fact, the quantum index is associated to a half of a separating real curve, while the other half has the opposite quantum index; for detailed definitions, see Section 3.2). Mikhalkin [19] showed that, for an appropriate kind of constraints, a Welschinger-type enumeration of real rational curves (cf. [25]) in a given divisor class and with a given quantum index produces an invariant and can be directly related to the numerator of a Block-Göttsche refined tropical invariant (represented as a fraction with the standard denominator).
@@
The main goal of the present paper is to introduce refined invariants of Mikhalkin’s type in the case of curves of genus $1$ and $2$. We follow the ideas of [23] and choose constraints so that every counted real curve of genus $g=1$ or $2$ appears to be a maximal one (i.e., it has $g+1$ global real branches), and hence is separating. More precisely, given a toric surface with the tautological real structure and a very ample divisor class, we fix maximally many real points in a generic Menelaus position (see Section 2.2) on the toric boundary of the positive quadrant, where genus $g$ curves from the given linear system must be tangent to toric divisors with prescribed even intersection multiplicities, and we fix $g$ more generic real points inside different non-positive quadrants as extra constraints (through which the curves under consideration should pass). There are finitely many real curves of genus $g$ matching the constraints and all these curves are separating. Their halves have quantum index, and we equip each curve with a certain Welschinger-type sign.
@@
For some toric surfaces (including the projective plane; in the case of genus $2$, in order to avoid cumbersome technicalities, this is the only toric surface considered in the paper), we prove that the signed enumeration of (halves of) real curves of genus $g=1$ or $2$ that match given constraints, belong to a given linear system, and have a prescribed quantum index does not depend on the choice of a (generic) position of the constraints; precise statements can be found in Section 3.3, Theorem 3.7, and Section 3.4, Theorem 3.9. The resulting invariants are said to be refined. In particular, we get new real enumerative invariants (without prescribing values for quantum index) in genus one and two. Tropical counterparts of the above refined invariants, their computation and comparison with the Block-Göttsche refined tropical invariants will be presented in the forthcoming papers [16, 17].
@@
The present paper is organized as follows. Section 2 contains preparatory material. In Section 3, we introduce refined invariants arising in enumeration of real curves of genus zero, one and two (the genus zero version slightly generalizes Mikhalkin’s refined rational invariants). Section 4 (respectively, 5, 6) is devoted to a proof of the invariance statement in genus zero (respectively, one, two).
@@ 2.1 Convex lattice polygons and real toric surfaces
Denote by ${\operatorname{Tor}}(P)$ the complex toric surface associated with $P$. Let ${\operatorname{Tor}}(P)^{\times}\simeq({\mathbb{C}}^{2})^{\times}$ (respectively, ${\operatorname{Tor}}(\partial P)$) be the dense orbit (respectively, the union of all toric divisors) of ${\operatorname{Tor}}(P)$. The toric surface ${\operatorname{Tor}}(P)$ has the tautological real structure, and the real part ${\operatorname{Tor}}_{\mathbb{R}}(P)$ of ${\operatorname{Tor}}(P)$ contains the positive quadrant ${\operatorname{Tor}}_{\mathbb{R}}(P)^{+}\simeq({\mathbb{R}}_{>0})^{2}$. The closure ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ of ${\operatorname{Tor}}_{\mathbb{R}}(P)^{+}$ (with respect to the Euclidean topology) is diffeomorphically taken onto $P$ by the moment map. We pull back the standard orientation and the metric of $P\subset{\mathbb{R}}^{2}$ to ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ and induce an orientation and a metric on the boundary $\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$; in particular, we get a cyclic order on the sides of $P$; we call this order positive. Denote by ${\mathcal{L}}_{P}$ the tautological line bundle over ${\operatorname{Tor}}(P)$; the global sections of ${\mathcal{L}}_{P}$ are spanned by the monomials $z^{\omega}$, $z=(z_{1},z_{2})$, $\omega\in P\cap{\mathbb{Z}}^{2}$.
@@ 2.2 Toric degree and Menelaus condition
For each $\sigma\in P^{1}_{\Delta}$, consider a sequence ${\boldsymbol{z}}^{\sigma}$ of $n^{\sigma}$ points in ${\operatorname{Tor}}(\sigma)^{\times}$, and denote by ${\boldsymbol{z}}$ the double index sequence (the upper index being $\sigma\in P^{1}_{\Delta}$ and the lower index being $i=1,\ldots,n^{\sigma}$) formed by the sequences ${\boldsymbol{z}}^{\sigma}$. We say that the sequence ${\boldsymbol{z}}$ satisfies the Menelaus $\Delta$-condition (cf. [19, Section 5.1]) if there exists a curve $C\in|{\mathcal{L}}_{P_{\Delta}}|$ such that $C$ does not contain toric divisors as components and, for each $\sigma\in P^{1}_{\Delta}$, the scheme-theoretic intersection of $C$ and ${\operatorname{Tor}}(\sigma)$ coincides with $\sum_{i=1}^{n^{\sigma}}2k^{\sigma}_{i}z^{\sigma}_{i}$.
@@
Note that the introduced above isomorphism ${\operatorname{Tor}}(\sigma)^{\times}\simeq{\mathbb{C}}^{\times}$ takes ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)^{+}$ onto ${\mathbb{R}}_{>0}\subset{\mathbb{C}}^{\times}$. For any $\tau>0$, denote by $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ the part of $\operatorname{Men}(\Delta)_{red}$ specified by the condition that, for each $\sigma\in P^{1}_{\Delta}$, all points of the subsequence ${\boldsymbol{w}}^{\sigma}$ belong to ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$ and lie at the distance $\geq\tau$ from the endpoints of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$.
@@
For a sufficiently small $\tau>0$, the (metric) closure $\overline{\operatorname{Men}}_{\mathbb{R}}^{\tau}(\Delta)\subset\prod_{\sigma\subset\partial P}({\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma))^{n^{\sigma}}$ is diffeomorphic to a convex polytope.
Proof. Since $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ is given by equation (1) with positive variables $\xi_{i}^{\sigma}$ satisfying restrictions of the form $0<{\operatorname{const}}_{1}\leq\xi_{i}^{\sigma}\leq{\operatorname{const}}_{2}<\infty$, the coordinate-wise logarithm takes $\overline{\operatorname{Men}}_{\mathbb{R}}^{\tau}(\Delta)$ onto a hyperplane section of a convex polytope. $\Box$
@@ 2.3 Curves on toric surfaces
Given a morphism ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ of a curve $\widehat{C}$ to ${\operatorname{Tor}}(P_{\Delta})$, we denote by
@@
  • ${\Phi}(\widehat{C})$ the reduced model of ${\Phi}_{*}(\widehat{C})$, where all components are taken with multiplicity one;
  • ${\Phi}^{*}(D)$, where $D\subset{\operatorname{Tor}}(P_{\Delta})$ is a divisor intersecting ${\Phi}(\widehat{C})$ in a finitely many points, the divisor on $\widehat{C}$ which is the pull-back of the scheme-theoretic intersection $D\cap{\Phi}_{*}(\widehat{C})$.
Denote by ${\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$ the space of isomorphism classes of maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$, where $\widehat{C}$ is a smooth connected curve of genus $g$ and ${\boldsymbol{p}}$ is a sequence of $m$ distinct points in $\widehat{C}$, such that ${\Phi}_{*}(\widehat{C})\in|{\mathcal{L}}_{P_{\Delta}}|$. Correspondingly, we denote by ¯$\overline{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$ the Kontsevich compactification of the above moduli space; this compactification is obtained by adding isomorphism classes of stable maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$, where $\widehat{C}$ is a connected nodal curve of arithmetic genus $g$ and ${\boldsymbol{p}}$ is a sequence of $m$ distinct points in ${\widehat{C}}\setminus{\operatorname{Sing}}\hskip 1.70709pt({\widehat{C}})$, such that ${\Phi}_{*}(\widehat{C})\in|{\mathcal{L}}_{P_{\Delta}}|$ (here ${\operatorname{Sing}}\hskip 1.70709pt({\widehat{C}})$ stands for the set of singularities of $\widehat{C}$). In what follows, we work with certain subspaces of these moduli spaces specified for genus $g=0$ in Section 3.2, for genus $g=1$ in Section 3.3, and for genus $2$ in Section 3.4). For any subset ${\mathcal{M}}\subset{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$, denote by $\overline{\mathcal{M}}$ the closure of $\mathcal{M}$ in the compactified moduli space ¯$\overline{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$.
We denote by $(C_{1}\cdot C_{2})_{z}$ the intersection multiplicity of curves $C_{1},C_{2}\subset{\operatorname{Tor}}(P_{\Delta})$ at a smooth point $z$ of ${\operatorname{Tor}}(P_{\Delta})$. By $C_{1}C_{2}$ we mean the total intersection multiplicity of the curves $C_{1},C_{2}\subset{\operatorname{Tor}}(P_{\Delta})$ provided that all intersection points of these curves are smooth in ${\operatorname{Tor}}(P_{\Delta})$. Given a local branch $B$ of a curve germ $(C,z)\subset{\operatorname{Tor}}(P_{\Delta})$ at a smooth point $z\in{\operatorname{Tor}}(P_{\Delta})$, we denote by ${\operatorname{ord}}B$ the intersection multiplicity of $B$ with a generic smooth curve through $z$.
@@
Suppose that the sequence ${\boldsymbol{w}}$ is in general position subject to the location with respect to the toric divisors as indicated above. Then, the space ${\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$ is finite. Moreover, for each element $\left[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})\right]\in{\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$, the map ${\Phi}$ takes $\widehat{C}$ birationally onto an immersed curve $C={\Phi}(\widehat{C})$ that, for each $\sigma\in P^{1}_{\Delta}$, intersects the toric divisor ${\operatorname{Tor}}(\sigma)$ at $n^{\sigma}$ distinct points and $C$ is smooth at each of these intersection points.
Proof. Consider the family ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$ formed by the elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}}_{\partial})\to{\operatorname{Tor}}(P_{\Delta})]$ having ${\Phi}$ birational onto its image and matching the condition (M2) above. We claim that

### v2

@@ Abstract
We introduce new invariants of a class of toric surfaces (including the projective plane) that arise from appropriate enumeration of real curves of genus one and two. These invariants admit a refinement similar to the one introduced by Grigory Mikhalkin in the rational case.
MSC-2010 classification: Primary 14N10, Secondary 14J26, 14P05
@@ 1 Introduction
Refined enumerative geometry, initiated in [1, 8], became one of the central topics in enumerative geometry with important links to closed and open Gromov-Witten invariants and to Donaldson-Thomas invariants. In a big part of known examples, refined invariants appear as one-parameter deformations of complex enumerative invariants (see, for example, [1, 5, 7, 8]). In his groundbreaking paper [24], G. Mikhalkin has proposed a refined invariant provided by enumeration of real rational curves and has related this invariant to the refined tropical invariants of F. Block and L. Göttsche [1]. Namely, he introduced an integer-valued quantum index for real algebraic curves in toric surfaces. The quantum index is defined in [24] for real curves satisfying the following assumptions: the curve has to intersect toric divisors only at real points and to be irreducible and separating; the latter condition means that, in the complex point set of the normalization of the curve, the complement of the real part is disconnected, i.e., formed by two halves exchanged by the complex conjugation (in fact, the quantum index is associated to a half of a separating real curve, while the other half has the opposite quantum index; for detailed definitions, see Section 3.2). Mikhalkin [24] showed that, for an appropriate kind of constraints, a Welschinger-type enumeration of real rational curves (cf. [32]) in a given divisor class and with a given quantum index produces an invariant which can be directly related to the numerator of a Block-Göttsche refined tropical invariant (represented as a fraction with the standard denominator).
@@
The main goal of the present paper is to introduce refined invariants of Mikhalkin’s type in the case of curves of genus $1$ and $2$. We follow the ideas of [29] and choose constraints so that every counted real curve of genus $g=1$ or $2$ appears to be a maximal one (i.e., it has $g+1$ global real branches), and hence is separating. More precisely, given a toric surface with the tautological real structure and a very ample divisor class, we fix maximally many real points in a generic Menelaus position (see Section 2.2) on the toric boundary of the positive quadrant, where genus $g$ curves from the given linear system must be tangent to toric divisors with prescribed even intersection multiplicities, and we fix $g$ more generic real points inside different non-positive quadrants as extra constraints (through which the curves under consideration should pass). There are finitely many real curves of genus $g$ matching the constraints, and all these curves are separating. Their halves have quantum index, and we equip each curve with a certain Welschinger-type sign. This gives rise to a signed enumeration of halves of real curves of genus $g=1$ and $2$ that match given constraints, belong to a given linear system, and have a prescribed quantum index $\kappa$ (we also treat the case $g=0$, slightly generalizing Mikhalkin’s enumeration to the situation with arbitrary even intersection multiplicities).
@@
We prove that, for some toric surfaces (including the projective plane; in the case of genus $2$, in order to avoid cumbersome technicalities, this is the only toric surface considered in the paper), the result $W_{g}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ of such enumeration, where the sequence $\Delta$ encodes the orders of tangencies and the sequence $\widehat{\boldsymbol{w}}$ encodes the position of the constraints, does not depend on the choice of a (generic) sequence $\widehat{\boldsymbol{w}}$; precise statements can be found in Section 3.2 (rational case, Theorem 3.2), Section 3.3 (elliptic case, Theorem 3.6), and Section 3.4 (genus two case, Theorem 3.8). The obtained invariants are said to be refined. In particular, we get new real enumerative invariants (without prescribing values for quantum index) in genus one and two. Tropical counterparts of the above refined invariants, their computation and comparison with the Block-Göttsche refined tropical invariants are presented in the forthcoming papers [19, 20].
The proofs of invariance are decomposed into four steps (these steps are common for all genera $g=0$, $1$, and $2$):
  • (1)
    identification of walls, i.e., loci of codimension one in the space of constraints, corresponding to degenerations in the set of counted curves,
  • (2)
    analysis of (local) deformations of germs of degenerate curves at their singular points,
  • (3)
    establishing transversality conditions that ensure the existence of one-dimensional families in the space of constraints transversally crossing the walls and realizing prescribed local deformations of degenerate counted curves,
  • (4)
    verification of the persistence of the count in all wall-crossing events.
In addition to the signed enumeration mentioned above, we suggest another choice for a Welschinger-type sign in the enumerative problems considered. This choice also yields an invariant enumeration. The proofs of the invariance for these two signed enumerations are literally the same, so we treat them together. It turns out that the results $W_{g}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ and ${\widetilde{W}}_{g}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ of two signed enumerations (see Section 3 for the notations) coincide in the case $g=0$ and in the case $g=1$ (and seem to coincide in the case $g=2$), up to a global sign depending only on $\Delta$ and $\kappa$. This statement is proved in the forthcoming paper [19] using the tropical calculation of these invariants. In fact, for the constraints that are sufficiently close to the tropical limit, the two introduced signs coincide (up to the above mentioned global sign) for individual curves under enumeration. Probably, it is also the case in general, without the additional assumption on the constraints.
@@
The paper is organized as follows. Section 2 contains a preparatory material. In Section 3, we introduce refined invariants arising in enumeration of real curves of genus zero, one and two (the genus zero version slightly generalizes Mikhalkin’s refined rational invariants [24]). Sections 4, 5, and 6 are devoted to the proofs of the invariance statements in genus zero, one, and two, respectively.
@@ 2.1 Convex lattice polygons and real toric surfaces
Denote by ${\operatorname{Tor}}(P)$ the complex toric surface associated with $P$. Let ${\operatorname{Tor}}(P)^{\times}\simeq({\mathbb{C}}^{*})^{2}$ (respectively, ${\operatorname{Tor}}(\partial P)$) be the dense orbit (respectively, the union of all toric divisors) of ${\operatorname{Tor}}(P)$. The toric surface ${\operatorname{Tor}}(P)$ has the tautological real structure, and the real part ${\operatorname{Tor}}_{\mathbb{R}}(P)$ of ${\operatorname{Tor}}(P)$ contains the positive quadrant ${\operatorname{Tor}}_{\mathbb{R}}(P)^{+}\simeq({\mathbb{R}}_{>0})^{2}$. The closure ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ of ${\operatorname{Tor}}_{\mathbb{R}}(P)^{+}$ (with respect to the Euclidean topology) is diffeomorphically taken onto $P$ by the moment map. We pull back the standard orientation and the metric of $P\subset{\mathbb{R}}^{2}$ to ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ and induce an orientation and a metric on the boundary $\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$ of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P)$; in particular, we get a cyclic order on the sides of $P$; we call this order positive. Denote by ${\mathcal{L}}_{P}$ the tautological line bundle over ${\operatorname{Tor}}(P)$; the global sections of ${\mathcal{L}}_{P}$ are spanned by the monomials $z^{\omega}$, $z=(z_{1},z_{2})$, $\omega\in P\cap{\mathbb{Z}}^{2}$.
@@ 2.2 Toric degree and Menelaus condition
For each $\sigma\in P^{1}_{\Delta}$, consider a sequence ${\boldsymbol{z}}^{\sigma}$ of $n^{\sigma}$ points in ${\operatorname{Tor}}(\sigma)^{\times}$, and denote by ${\boldsymbol{z}}$ the double index sequence (the upper index being $\sigma\in P^{1}_{\Delta}$ and the lower index being $i=1,\ldots,n^{\sigma}$) formed by the sequences ${\boldsymbol{z}}^{\sigma}$. We say that the sequence ${\boldsymbol{z}}$ satisfies the Menelaus $\Delta$-condition (cf. [24, Section 5.1]) if there exists a curve $C\in|{\mathcal{L}}_{P_{\Delta}}|$ such that $C$ does not contain toric divisors as components and, for each $\sigma\in P^{1}_{\Delta}$, the scheme-theoretic intersection of $C$ and ${\operatorname{Tor}}(\sigma)$ coincides with $\sum_{i=1}^{n^{\sigma}}2k^{\sigma}_{i}z^{\sigma}_{i}$.
@@
Note that the introduced above isomorphism ${\operatorname{Tor}}(\sigma)^{\times}\simeq{\mathbb{C}}^{\times}$ takes ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)^{+}$ onto ${\mathbb{R}}_{>0}\subset{\mathbb{C}}^{\times}$. For $\rho>0$, denote by $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ the part of $\operatorname{Men}(\Delta)_{red}$ specified by the condition that, for each $\sigma\in P^{1}_{\Delta}$, all points of the subsequence ${\boldsymbol{w}}^{\sigma}$ belong to ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$ and lie at the distance $\geq\rho$ (with respect to the metric fixed in Section 2.1) from the endpoints of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$. Here, we always assume that, in each interval ${\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma)$, the complement to $\rho$-neighborhoods of the endpoints is non-empty.
@@
The (metric) closure $\overline{\operatorname{Men}}_{\mathbb{R}}^{\rho}(\Delta)\subset\prod_{\sigma\subset\partial P}({\operatorname{Tor}}_{\mathbb{R}}^{+}(\sigma))^{n^{\sigma}}$ is diffeomorphic to a convex polytope.
Proof. Since $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ is given by equation (1) with positive variables $\xi_{i}^{\sigma}$ satisfying restrictions of the form $0<{\operatorname{const}}_{1}\leq\xi_{i}^{\sigma}\leq{\operatorname{const}}_{2}<\infty$, the coordinate-wise logarithm takes $\overline{\operatorname{Men}}_{\mathbb{R}}^{\rho}(\Delta)$ onto a hyperplane section of a convex polytope. $\Box$
2.3 Deformations of plane curve germs
Here we introduce notations and recall some known facts on deformations of plane curve germs.
Let $(C,z)\subset({\mathbb{C}}^{2},z)$ be the germ of a reduced curve $C$ with an isolated singular point $z$, and let $L\supset\{z\}$ be a smooth curve intersecting $(C,z)$ only at $z$. The space $\mathrm{Vers}(C,z)$, defined as the germ at zero of ${\mathcal{O}}_{C,z}/{\mathfrak{m}}_{C,z}^{s}$, $s\gg 0$, where ${\mathfrak{m}}_{C,z}$ is the maximal ideal in the local ring ${\mathcal{O}}_{C,z}$, serves as the base of a versal deformation of the germ $(C,z)$. Denote by $Q_{1},...,Q_{r}$ the irreducible components of $(C,z)$, i.e., the local branches of $C$ centered at $z$, and consider the following loci in $\mathrm{Vers}(C,z)$.
  1. (a)
    The equisingular locus $ES(C,z)\subset\mathrm{Vers}(C,z)$ parameterizes curve germs with a singular point topologically equivalent to $(C,z)$.
    It is smooth (see [4, Item (ii) on page 435]), and by [12, Theorem 1 and formula (19)], its tangent space $T(ES(C,z))$ is contained in
    $\left\{\varphi\in\mathrm{Vers}(C,z)\ \Big{|}\ \begin{array}[]{c}{{\operatorname{ord}}\varphi\big{|}_{Q_{i}}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {\qquad\qquad+{\operatorname{mult}}Q_{i}-1,\quad i=1,...,r}\end{array}\right\},$
    where $(Q_{i}\cdot Q_{j})$ stands for the intersection multiplicity of the local branches $Q_{i}$ and $Q_{j}$, the $\delta$-invariant $\delta(Q_{i})$ denotes the maximal number of nodes that may appear in a deformation of the local branch $Q_{i}$, the multiplicity of $Q_{i}$ at $z$ is denoted by ${\operatorname{mult}}Q_{i}$, and at last, ${\operatorname{ord}}\varphi\big{|}_{Q_{i}}$ is the vanishing order of $\varphi(x(t),y(t))$, where $(x(t),y(t))$ is a parameterization of $Q_{i}$.
  2. (b)
    The equisingular loci $ES_{L}(C,z)\subset ES(C,z)$ and $ES^{fix}_{L}(C,z)\subset ES_{L}(C,z)$ additionally constrained by the requirement to intersect $L$ at a point close to $z$ with multiplicity $(C\cdot L)_{z}$, or to intersect $L$ at $z$ with multiplicity $(C\cdot L)_{z}$, respectively. Both loci are smooth, and by [12, Theorem 2 and formulas (21), (22)], their tangent spaces satisfy
    $T(ES_{L}(C,z))\subset\left\{\varphi\in\mathrm{Vers}(C,z)\ \Big{|}\ \begin{array}[]{c}{{\operatorname{ord}}\varphi\big{|}_{Q_{i}}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {\qquad\qquad+(Q_{i}\cdot L)-1,\quad i=1,...,r}\end{array}\right\},$
    $T(ES^{fix}_{L}(C,z))\subset\left\{\varphi\in\mathrm{Vers}(C,z)\ \Big{|}\ \begin{array}[]{c}{{\operatorname{ord}}\varphi\big{|}_{Q_{i}}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {+{\operatorname{mult}}Q_{i}+(Q_{i}\cdot L)-1,\quad i=1,\ldots,r}\end{array}\right\}.$
  3. (c)
    The equigeneric locus $EG(C,z)\subset\mathrm{Vers}(C,z)$ parameterizing equigeneric (or $\delta={\operatorname{const}}$) deformations. It is smooth if and only if all local branches $Q_{1}$, …, $Q_{r}$ are smooth. In any case, its tangent cone is the linear space
    $J^{cond}(C,z)/{\mathfrak{m}}_{C,z}^{s}=\left\{\varphi\in\mathrm{Vers}(C,z)\ \Big{|}\ \begin{array}[]{c}{{\operatorname{ord}}\varphi\big{|}_{Q_{i}}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {i=1,\ldots,r}\end{array}\right\}$
    of codimension ${\operatorname{codim}}EG(C,z)=\delta(C,z)$, where
    $J^{cond}(C,z)=\operatorname{Ann}({\Phi}_{*}{\mathcal{O}}_{\widehat{C}}/{\mathcal{O}}_{C})\subset{\mathcal{O}}_{C,z}$
    is the conductor ideal (see details in [4, Item (iii) on page 435 and Theorem 4.15]), and $\delta(C,z)=\dim{\mathcal{O}}_{C,z}/J^{cond}(C,z)$.
  4. (d)
    The equigeneric loci $EG_{L}(C,z)\subset EG(C,z)$ and $EG_{L}^{fix}(C,z)\subset EG_{L}(C,z)$ additionally constrained by the requirement to intersect $L$ at a point close to $z$ with multiplicity $(C\cdot L)_{z}$, or to intersect $L$ at $z$ with multiplicity $(C\cdot L)_{z}$, respectively. If $(C,z)$ is irreducible, i.e., $r=1$, then these loci are smooth [30, Lemma 3(1)], and their tangent spaces satisfy
    $T(EG_{L}(C,z))=\left\{\varphi\in\mathrm{Vers}(C,z)\ \Big{|}\ {\operatorname{ord}}\varphi\big{|}_{(C,z)}\geq 2\delta(C,z)+(C\cdot L)_{z}-1\right\},$
    $T(EG^{fix}_{L}(C,z))=\left\{\varphi\in\mathrm{Vers}(C,z)\ \Big{|}\ {\operatorname{ord}}\varphi\big{|}_{(C,z)}\geq 2\delta(C,z)+(C\cdot L)_{z}\right\}.$
@@ 2.4 Curves on toric surfaces
Given a morphism ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ of a smooth curve $\widehat{C}$ to ${\operatorname{Tor}}(P_{\Delta})$, we denote by
@@
  • $C:={\Phi}(\widehat{C})$ the reduced model of ${\Phi}_{*}(\widehat{C})$, where all components are taken with multiplicity one; we split ${\Phi}$ into ${\Phi}:\widehat{C}\twoheadrightarrow C\overset{{\operatorname{in}}}{\hookrightarrow}{\operatorname{Tor}}(P_{\Delta})$;
  • ${\Phi}^{*}(D)$, where $D\subset{\operatorname{Tor}}(P_{\Delta})$ is a divisor intersecting $C$ in finitely many points, the divisor on $\widehat{C}$ which is the pull-back of the scheme-theoretic intersection $D\cap{\Phi}_{*}(\widehat{C})$.
Denote by ${\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$ the moduli space of stable maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$, where $\widehat{C}$ is a smooth connected curve of genus $g$ and ${\boldsymbol{p}}$ is a sequence of $m$ distinct points in $\widehat{C}$, such that ${\Phi}_{*}(\widehat{C})\in|{\mathcal{L}}_{P_{\Delta}}|$. Correspondingly, we denote by ¯$\overline{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$ the Kontsevich compactification of the above moduli space; this compactification is obtained by adding isomorphism classes of stable maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$, where $\widehat{C}$ is a connected nodal curve of arithmetic genus $g$ and ${\boldsymbol{p}}$ is a sequence of $m$ distinct points in ${\widehat{C}}\setminus{\operatorname{Sing}}\hskip 1.70709pt({\widehat{C}})$, such that ${\Phi}_{*}(\widehat{C})\in|{\mathcal{L}}_{P_{\Delta}}|$ (here ${\operatorname{Sing}}\hskip 1.70709pt({\widehat{C}})$ stands for the set of singularities of $\widehat{C}$). In what follows, we work with certain subspaces of these moduli spaces specified for genus $g=0$ in Section 3.2, for genus $g=1$ in Section 3.3, and for genus $2$ in Section 3.4. For any subset ${\mathcal{M}}\subset{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$, denote by $\overline{\mathcal{M}}$ the closure of $\mathcal{M}$ in the compactified moduli space ¯$\overline{\mathcal{M}}_{g,m}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$.
We denote by $(C_{1}\cdot C_{2})_{z}$ the intersection multiplicity of curves $C_{1},C_{2}\subset{\operatorname{Tor}}(P_{\Delta})$ at a smooth point $z$ of ${\operatorname{Tor}}(P_{\Delta})$. By $C_{1}C_{2}$ we mean the total intersection multiplicity of the curves $C_{1},C_{2}\subset{\operatorname{Tor}}(P_{\Delta})$ provided that all intersection points of these curves are smooth in ${\operatorname{Tor}}(P_{\Delta})$.
@@
In case of ${\Phi}:\widehat{C}\to C\hookrightarrow{\operatorname{Tor}}(P_{\Delta})$ birational onto its image $C$, but not necessarily immersion, we consider deformations realizing specific local deformations of $C$ at its singular points (see details in Section 2.3).
@@
Suppose that the sequence ${\boldsymbol{w}}$ is in general position subject to the location with respect to the toric divisors as indicated above. Then, the space ${\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$ is finite. Moreover, for each element $\left[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})\right]\in{\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$, the map ${\Phi}$ takes $\widehat{C}$ birationally onto an immersed curve $C={\Phi}(\widehat{C})$ that, for each $\sigma\in P^{1}_{\Delta}$, intersects the toric divisor ${\operatorname{Tor}}(\sigma)$ at $n^{\sigma}$ distinct points and $C$ is smooth at each of these intersection points. Furthermore, ${\Phi}({\boldsymbol{p}})\cap{\operatorname{Sing}}\hskip 1.70709pt(C)=\emptyset$.
Proof. Consider the family
${\mathcal{M}}^{\prime}_{g,n}(\Delta)=\big{\{}[{\Phi}:(\widehat{C},{\boldsymbol{p}}_{\partial})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{g,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$
such that
${\Phi}:\widehat{C}\to C={\Phi}(\widehat{C})\ \text{is birational},\qquad\qquad$
[…]
We claim that